# RF4D:Neural Radar Fields for Novel View Synthesis in Outdoor Dynamic Scenes

### v2

@@ Abstract
Neural fields (NFs) have demonstrated remarkable performance in scene reconstruction, powering various tasks such as novel view synthesis. However, existing NF methods relying on RGB or LiDAR inputs often exhibit severe fragility to adverse weather, particularly when applied in outdoor scenarios like autonomous driving. In contrast, millimeter-wave radar is inherently robust to environmental changes, while unfortunately, its integration with NFs remains largely underexplored. Besides, as outdoor driving scenarios frequently involve moving objects, making spatiotemporal modeling essential for temporally consistent novel view synthesis. To this end, we introduce RF4D, a radar-based neural field framework specifically designed for novel view synthesis in outdoor dynamic scenes. RF4D explicitly incorporates temporal information into its representation, significantly enhancing its capability to model moving objects. We further introduce a feature-level flow module that predicts latent temporal offsets between adjacent frames, enforcing temporal coherence in dynamic scene modeling. Moreover, we propose a radar-specific power rendering formulation closely aligned with radar sensing physics, improving synthesis accuracy and interpretability. Extensive experiments on public radar datasets demonstrate the superior performance of RF4D in terms of radar measurement synthesis quality and occupancy estimation accuracy, achieving especially pronounced improvements in dynamic outdoor scenarios.
@@ 1 Introduction
Dynamic scene reconstruction play a vital role in various fields such as augmented/virtual reality (AR/VR) [1], robotics [2] and autonomous driving [3]. Recent advances in computer vision have enabled high-fidelity 3D scene reconstruction and novel view synthesis (NVS), which support downstream tasks including simulation [4], motion planning [5, 6], and scene understanding [7, 8]. For instance, dynamic scene reconstruction enables closed-loop simulations that are essential for training and evaluating end-to-end planning algorithms under diverse and complex scenarios [9].
While NVS has been extensively explored using camera [10, 11, 12]and LiDAR data [13, 14], radar remains largely underutilized despite its widespread deployment in autonomous driving systems. Radar sensors offer key advantages such as robustness to adverse weather and low light, long-range sensing, and low cost [15]. Like LiDAR and RGB sensors, radar also provides partial observations that vary across viewpoints, making it a suitable candidate for scene reconstruction. However, radar presents unique challenges due to its low spatial resolution, sparsity, internal noise, and multipath effects [16]. Moreover, outdoor driving scenarios frequently involve moving objects, further complicating dynamic scene reconstruction from radar measurements.
Traditional radar-based scene reconstruction methods process individual measurements into sparse point clouds, which are then aggregated across frames and transformed into a global coordinate system. To synthesize radar measurements from novel viewpoints, prior works employ simulation techniques such as ray tracing and radio frequency (RF) propagation modeling [17, 18]. However, these approaches rely heavily on access to radar hardware and waveform specifications, making them difficult to generalize across different radars. Furthermore, they are primarily limited to static scenes and often fail to accurately capture view-dependent radar cross section (RCS) characteristics of real-world objects.
Neural fields can implicitly represent scenes and have driven significant advances in scene reconstruction and novel view synthesis. Their representative work, Neural Radiance Fields (NeRF) [19], combines the neural fields with optical volume rendering techniques, delivering remarkable performance for RGB images. Recent work has explored the use of NFs for radar-based novel view synthesis. Radar Fields [20], in particular, introduces bypassing optical volume rendering by using a physics-motivated forward model to predict the received radar power at each location in the scene. However, we observe that their forward model leads to occupancy-reflectance contradiction: high predicted occupancy often leads to suppressed reflectance (as shown in Figure 1b). These results violate the expected relationship between occupancy and reflectance, where high occupancy should typically correspond to stronger reflections. Moreover, these methods focus on static scenes, ignoring the dynamic nature of outdoor driving scenarios, and often fail to handle moving targets. Figure 1a shows one example in which a moving vehicle disappears in the novel view. In addition, their methods rely on an external occupancy estimator for supervision, making it sensitive to estimator quality and breaking the end-to-end training pipeline.
To tackle the aforementioned limitations, we propose RF4D, a radar-based neural field framework for NVS in outdoor dynamic scenes. Specifically, we represent the dynamic scenes using neural radar fields by taking both position and time as input, and predict two radar-specific characteristics at each point: occupancy, indicating whether the position is occupied or empty, and radar cross section (RCS), describing how detectable an object is, if present. Building upon this framework, we further introduce a feature-level flow module that predicts residual temporal offsets in latent space to encourage stable and consistent occupancy predictions across frames. To resolve the occupancy-reflectance contradiction, we revisit the radar signal formation model and propose a rigorous power rendering formulation to estimate the received power at each position using the predicted occupancy and RCS values. Our formulation reserves physically faithful relationships between RCS and scene occupancy and removes the need for external occupancy supervision. As shown in Figure 1b, our method produces occupancy and RCS predictions that are consistent with radar physics where high occupancy corresponds to strong reflectance, rather than the contradictory behavior observed in Radar Fields. In summary, our contributions can be summarized as follows:
  • We propose RF4D, a radar-based neural field framework for novel view synthesis in outdoor dynamic scenes, by representing the scene as a function of both position and time. Besides, we introduce a feature-level flow module to encourage stable occupancy predictions across adjacent time frames by learning residual temporal offsets in latent space.
  • We design a radar-specific power rendering formulation that estimates the received power using predicted occupancy and RCS values, which preserves physically faithful relationships and eliminates the need for external occupancy supervision.
  • We demonstrate the effectiveness of the proposed method with two public radar datasets compared with SOTA, in terms of synthesizing quality and occupancy estimation accuracy.
@@
Figure 1: (a) Comparison of radar view synthesis results for a dynamic scene with a moving vehicle (red box). Our method successfully renders the moving object, while Radar Fields [20] fails to recover it. (b) Predicted occupancy and reflectance (Radar Fields [20]) versus occupancy and RCS (ours). Our predictions follow radar physics, where high occupancy corresponds to strong RCS, while Radar Fields [20] fails to maintain such consistency between occupancy and reflectance.
@@ 2.1 Radar simulation
Radar simulation aim to generate realistic radar measurements for scenes. Traditional radar simulation methods rely on physical modeling of radar sensors and the environments through ray tracing and Radio Frequency propagation modeling [17, 21, 22, 23, 24, 18, 25], but they require detailed knowledge of radar hardware and do not scale across sensors. Generative models, such as generative adversarial networks (GAN) or variational autoencoders (VAE), have been used to learn radar data distributions from real measurements, but they depend on large annotated datasets and sensor-specific tuning [26, 27, 28]. RadSimReal [29] introduces a learning-based radar simulator that synthesizes realistic radar signals from 3D ground-truth scenes, without requiring detailed radar hardware modeling. Recent efforts have adapted neural fields to radar by modeling occupancy and reflectance. DART [30] represents range-Doppler measurements using neural fields collected with handheld radar. Radar Fields [20] targets range-azimuth representations using ego-motion radar, relying on external occupancy estimators [31] for regularization. These works demonstrate the feasibility of NF-based radar synthesis but they are limited to static scenes. We extend this line of research by enabling dynamic scene modeling without additional occupancy supervision.
@@ 4 Method
Figure 2: Overview of the proposed RF4D framework. Given a 3D query point $(x,y,z)$ at time $t$ and view direction $\mathbf{d}$, RF4D first predicts two radar-specific physical quantities: occupancy $\alpha$ and radar cross section (RCS) $\sigma$, using neural radar fields. The predicted occupancy reflects whether the point is physically occupied, while the RCS describes its reflectivity. Then these quantities are combined via our proposed radar-specific power rendering to estimate the received radar power. During training, besides the supervision of ground truth radar measurements, the feature-level flow module promotes temporal consistency of occupancy by modeling latent feature changes across adjacent frames.
@@ 4.1 Problem Formulation
In dynamic driving scenarios, we are given a sequence of radar measurements $S=\{S_{0},S_{1},...,S_{n-1}\}$, where $S_{i}\in\mathbb{R}^{N_{\theta}\times N_{\delta}}$, along with the corresponding radar poses $H_{s}=\{H_{0},H_{1},...,H_{n-1}\}$, where $H_{i}\in\mathrm{SE}(3)$, and timestamps $T_{s}=\{t_{0},t_{1},...,t_{n-1}\}$, where $t_{i}\in\mathbb{R}$. Here, $N_{\theta}$ denotes the number of azimuth directions (beams), and $N_{\delta}$ is the number of range bins per beam. Each bin records a 1D reflected power value measured at corresponding azimuth and range.
@@ 4.2 RF4D
The latent feature $\chi$ is then passed to $f_{\alpha}$, an MLP that predicts occupancy using a Gumbel-Sigmoid activation [48] at the output. Note that Gumbel-Sigmoid is to encourage the predicted $\alpha$ to approximate binary values, functioning as a soft binary mask aligned with physical interpretation of occupancy: 0 implies empty and 1 implies occupied. Due to that RCS also depends on the incident angle of radar waveform (view direction), to predict RCS $\sigma$, we concatenate $\chi$ with the encoded direction $\mathbf{d}$ and pass them to $f_{\sigma}$. Here we use spherical harmonics $\mathcal{S}$ [49] to encode $\mathbf{d}$. Finally, we obtain the two predicted radar-specific characteristics:
@@
To ensure temporal consistency in occupancy predictions across frames, we introduce a feature-level flow module inspired by concepts in optical and scene flow estimation. This module captures the temporal dynamics of the scene by modling the evolution of latent features over time. Given a latent feature representation $\chi$ at time $t$, we aim to predict the occupancy at neighboring time frames $t-\Delta t$ and $t+\Delta t$. To achieve this, we concatenate the time offset $\Delta t$ with the latent feature $\chi$ and input this into an MLP $f_{\Delta\chi}$ to estimate the temporal offset $\Delta\chi$. This offset approximates how the feature at the current point changes across time:
$\chi^{\prime}=\chi+f_{\Delta\chi}(\chi,\Delta t).$ (4)
We then update the original feature by adding the predicted offset: $\chi^{\prime}=\chi+\Delta\chi$. This updated feature $\chi^{\prime}$ is passed through the occupancy prediction MLP $f_{\alpha}$ to obtain the occupancy estimates at the neighboring time frames:
$\alpha^{t-\Delta t}=f_{\alpha}(\chi+f_{\Delta\chi}(\chi,\Delta t)),$ (5)
$\alpha^{t+\Delta t}=f_{\alpha}(\chi+f_{\Delta\chi}(\chi,-\Delta t)).$
We then enforce temporal consistency by regularizing the occupancy predictions across time. This encourages the neural radar field to learn temporally coherent representations of occupancy, especially in dynamic scenes. Conceptually, this mimics the role of optical or scene flow: tracking how point-wise features change over time to maintain stable and geometry-aware predictions.
@@
Unlike volume rendering used in NeRF, synthesizing radar measurements requires incorporating radar sensing physics. In particular, the received radar power is governed by the physics model in Equation 1, where power depends on the RCS $\sigma$ of the target and its distance $\delta$ from the sensor. To simplify the model, we discard constant terms such as transmitted power and antenna gain, which are fixed and absorbed during training. Additionally, since radar measurements are typically provided in decibel (dB) scale, we apply a base-10 logarithmic transformation. Importantly, radar reflections only occur when a target is physically present. To account for this, we use the predicted occupancy $\alpha\in[0,1]$ as a soft gating term to modulate the power response. Combining these considerations, the radar-specific power rendering is defined as:
@@
This formulation respects radar sensing physics and enables consistent predictions between radar measurement and occupancy.

### v3

@@ Abstract
Neural fields (NFs) have achieved remarkable success in scene reconstruction and novel view synthesis. However, existing NF approaches that rely on RGB or LiDAR inputs often struggle under adverse weather conditions, limiting their robustness in real-world outdoor environments such as autonomous driving. In contrast, millimeter-wave radar is inherently resilient to environmental variations, yet its integration with NFs remains largely underexplored. Moreover, outdoor driving scenes frequently involve dynamic objects, making spatiotemporal modeling crucial for temporally consistent novel view synthesis. To address these challenges, we present RF4D, a radar-based neural field framework tailored for novel view synthesis in outdoor dynamic scenes. RF4D explicitly incorporates temporal information into its representation, enabling more accurate modeling of object motion. A dedicated scene flow module further predicts motion offsets between adjacent frames, enforcing temporal occupancy coherence during dynamic scene reconstruction. Moreover, we propose a radar-specific power rendering formulation grounded in radar sensing physics, improving both synthesis accuracy and interpretability. Extensive experiments on public radar datasets demonstrate that RF4D substantially outperforms existing methods in radar measurement synthesis and occupancy estimation accuracy, with particularly strong gains in dynamic outdoor environments. Code is available at https://zhan0618.github.io/RF4D.
@@ 1 Introduction
Figure 1: Comparison of radar view synthesis for a dynamic scene with a moving vehicle (orange box). RF4D successfully renders the moving object, whereas Radar Fields [4] fails to recover it.
Dynamic scene reconstruction plays a crucial role in numerous applications, including augmented and virtual reality (AR/VR) [9], robotics [70] and autonomous driving [54]. Recent advances in computer vision [27, 60, 34, 57, 20, 19, 58, 17, 18, 51] have enabled high-fidelity 3D scene reconstruction and novel view synthesis (NVS), which in turn support downstream tasks such as simulation [66], motion planning [23, 24], and scene understanding [46, 36]. For instance, dynamic scene reconstruction enables closed-loop simulations for training and evaluating end-to-end planning algorithms under diverse, real-world conditions [68]. While NVS has been widely studied using camera [39, 55, 61] and LiDAR data [52, 69], radar sensing remains relatively underexplored despite its extensive use in autonomous driving. Radar offers unique advantages, such as robustness to adverse weather and low light, long-range perception, and cost efficiency [56, 14]. Similar to LiDAR and RGB sensors, radar also provides partial and viewpoint-dependent observations, making it a natural candidate for scene reconstruction. However, radar data is particularly challenging due to its low spatial resolution, sparsity, internal noise, and multipath effects [45, 13, 15]. The presence of moving objects in outdoor driving scenes further complicates the reconstruction process.
Traditional radar-based reconstruction methods typically convert individual radar measurements into sparse point clouds, aggregate them across frames, and align them in a global coordinate system. To enable novel view synthesis, prior works often rely on simulation-based techniques such as ray tracing or radio frequency (RF) propagation modeling [1, 48]. However, these approaches rely heavily on hardware-specific radar parameters and waveform configurations, limiting their generalizability [3]. Moreover, they are mostly constrained to static scenes and struggle to capture view-dependent radar cross section (RCS) variations observed in real-world dynamic environments.
Neural fields (NFs) have recently emerged as a powerful paradigm for implicit scene representation, enabling breakthroughs in both reconstruction and NVS. The seminal Neural Radiance Fields (NeRF) [35] integrates NFs with optical volume rendering, achieving remarkable results for RGB-based reconstruction. Extending this idea, Radar Fields [4] introduced a radar-based NF model that replaces optical rendering with a physics-driven forward model to predict received radar power. However, they are limited to static scenes, leading to failure cases in dynamic environments (e.g., disappearing moving vehicles, as shown in Figure 1). Furthermore, we observe a fundamental issue in their design, i.e., an occupancy–reflectance contradiction, where regions with high predicted occupancy often exhibit suppressed reflectance, as illustrated in Figure 2. This behavior violates the physical expectation that occupied regions should produce stronger reflections. These models also rely on external occupancy estimators for supervision, introducing additional dependencies.
To address these limitations, we propose RF4D, a radar-based neural field framework designed for NVS in outdoor dynamic scenes. Specifically, RF4D represents each scene as a spatiotemporal NF parameterized by both position and time, and predicts two radar-specific properties at each spatial point: (1) occupancy, indicating whether the point is occupied, and (2) RCS, describing the reflectivity of the occupied region. To ensure temporal consistency, we introduce a scene flow module that predicts motion offsets, promoting stable and coherent occupancy predictions across adjacent frames. To resolve the occupancy–reflectance contradiction, we further propose a radar-specific power rendering formulation grounded in radar sensing physics, which estimates the received radar power based on the predicted occupancy and RCS. This formulation maintains a physically faithful relationship between occupancy and reflectance, while eliminating the need for external occupancy supervision. As illustrated in Figure 2, RF4D produces predictions consistent with radar physics, i.e., high occupancy corresponds to strong reflectance, unlike the contradictory behavior seen in prior work.
In summary, our contributions are three fold:
  • We propose RF4D, a radar-based neural field framework for novel view synthesis in dynamic outdoor scenes, representing the scene as a continuous function of both spatial and temporal dimensions. A scene flow module promotes temporal coherence by learning motion offsets across frames.
  • We design a radar-specific power rendering formulation that jointly leverages predicted occupancy and RCS, preserving physically consistent relationships and removing the dependence on external occupancy supervision.
  • We conduct extensive experiments on two public radar datasets, demonstrating that RF4D achieves state-of-the-art performance in radar measurement synthesis and occupancy estimation, particularly in dynamic scenarios.
Figure 2: Predicted occupancy and reflectance from Radar Fields [4] versus occupancy and radar cross-section (RCS) from RF4D. Our predictions follow radar physics, where high occupancy corresponds to strong RCS, while Radar Fields lacks such consistency.
Figure 3: Overview of the proposed RF4D framework. Given a 3D query point $x$ at time $t$ and view direction $\mathbf{d}$, RF4D first predicts two radar-specific physical quantities: occupancy $\alpha$ and radar cross-section (RCS) $\sigma$, using neural radar fields. The occupancy $\alpha$ indicates whether the point is physically occupied, and the RCS $\sigma$ represents its reflectivity. These quantities are combined through the radar-specific power rendering to estimate the received radar power. During training, the rendered power is supervised by ground-truth radar measurements, and the scene flow module enforces temporal consistency by predicting motion offsets and warping points to adjacent frames to regularize occupancy over time.
@@ 2.1 Radar simulation
Radar simulation aims to generate realistic radar measurements for scenes. Traditional radar simulation methods rely on physical modeling of radar sensors and the environments through ray tracing and Radio Frequency propagation modeling [1, 11, 21, 22, 47, 48, 53], but they require detailed knowledge of radar hardware and do not scale across sensors. Generative models, such as generative adversarial networks (GAN) or variational autoencoders (VAE), have been used to learn radar data distributions from real measurements, but they depend on large annotated datasets and sensor-specific tuning [64, 59, 65]. RadSimReal [3] introduces a learning-based radar simulator that synthesizes realistic radar signals from 3D ground-truth scenes, without requiring detailed radar hardware modeling. Recent efforts have adapted neural fields to radar by modeling occupancy and reflectance. NeuRadar [44] jointly models camera, LiDAR, and radar, and uses camera/LiDAR cues to guide radar point cloud rendering. DART [25] represents range-Doppler measurements using neural fields collected with handheld radar. Radar Fields [4] targets range-azimuth representations using ego-motion radar, relying on external occupancy estimators [62] for regularization. RadarSplat [29] advances detailed physical modeling of radar to synthesize realistic measurements, including multipath and noise. This direction is orthogonal to our approach and is broadly beneficial to radar NVS. Overall, existing methods address static scenes only and do not model dynamics.
@@ 3 Preliminary
Table 1: Quantitative comparison across four different driving scenarios from the RobotCar dataset [2]. The best result and the runner-up are highlighted in bold and underline, respectively.
Scene 1 Scene 2 Scene 3 Scene 4
Method SSIM↑ PSNR↑ CD↓ RCD↓ SSIM↑ PSNR↑ CD↓ RCD↓ SSIM↑ PSNR↑ CD↓ RCD↓ SSIM↑ PSNR↑ CD↓ RCD↓
D-NeRF [43] 0.1270 20.80 80.9668 0.1046 0.1363 20.29 96.7480 0.1114 0.1620 21.27 23.8396 0.1198 0.0867 17.58 69.4628 0.1710
Hexplane [7] 0.2674 21.12 78.9851 0.0879 0.2980 20.21 84.3646 0.2238 0.3909 22.09 119.0078 0.2068 0.2267 20.25 58.8482 0.2207
RadarFields [4] 0.3372 23.36 18.1915 0.0300 0.2997 22.24 16.1489 0.0149 0.3498 24.02 9.5357 0.0281 0.2829 21.95 20.9821 0.0219
Ours 0.6103 23.38 7.3449 0.0190 0.5350 22.16 8.0909 0.0125 0.6258 23.76 3.2896 0.0097 0.4971 22.20 18.7607 0.0147
Table 2: Quantitative comparison across four driving scenarios from the Boreas dataset [6]. The best result and the runner-up are highlighted in bold and underline, respectively. CD and RCD are not reported for the snow scene due to unreliable ground-truth LiDAR geometry under heavy snowfall.
Sun Snow Rain Static
Method SSIM↑ PSNR↑ CD↓ RCD↓ SSIM↑ PSNR↑ CD↓ RCD↓ SSIM↑ PSNR↑ CD↓ RCD↓ SSIM↑ PSNR↑ CD↓ RCD↓
D-NeRF [43] 0.0923 12.33 144.7542 1.4754 0.1676 20.48 - - 0.0991 13.31 435.1305 1.1928 0.1483 19.65 46.1635 0.2127
Hexplane [7] 0.2892 22.90 28.2160 0.0911 0.2643 23.61 - - 0.2946 23.28 38.5217 0.2149 0.2388 24.23 46.1504 0.2114
RadarFields [4] 0.3641 25.39 10.6385 0.0060 0.4823 27.36 - - 0.3905 25.80 23.8941 0.0312 0.4331 26.96 14.1116 0.0069
Ours 0.7001 26.65 5.4482 0.0033 0.7946 27.98 - - 0.6635 26.21 10.4558 0.0051 0.7724 27.55 8.0739 0.0040
Scene 1 Scene 2 Scene 3 Scene 4
$\mathcal{L}_{p}$ $\mathcal{L}_{oc}$ $\mathcal{L}_{m}$ PSNR↑ SSIM↑ CD↓ RCD↓ PSNR↑ SSIM↑ CD↓ RCD↓ PSNR↑ SSIM↑ CD↓ RCD↓ PSNR↑ SSIM↑ CD↓ RCD↓
21.90 0.3184 80.9668 0.1046 20.35 0.2616 96.7480 0.1114 23.80 0.4161 23.8396 0.1198 22.55 0.3586 45.3407 0.1174
23.23 0.5662 9.2251 0.0200 22.10 0.5315 11.8233 0.0145 23.56 0.6240 6.5126 0.0108 22.10 0.5152 23.8087 0.0176
22.99 0.6167 7.5713 0.0150 22.08 0.5269 10.4477 0.0137 23.33 0.6182 5.1028 0.0098 21.58 0.4933 19.5720 0.0150
23.38 0.6103 7.3449 0.0190 22.16 0.5350 8.0909 0.0125 23.76 0.6258 3.2896 0.0097 22.20 0.4971 18.7607 0.0147
Table 4: Ablation study for the Boreas dataset [6]. The best result and the runner-up are highlighted in bold and underline, respectively. CD and RCD are not reported for the snow scene due to unreliable ground-truth LiDAR geometry under heavy snowfall.
Sun Snow Rain Static
$\mathcal{L}_{p}$ $\mathcal{L}_{oc}$ $\mathcal{L}_{m}$ PSNR ↑ SSIM ↑ CD↓ RCD↓ PSNR ↑ SSIM↑ CD↓ RCD↓ PSNR↑ SSIM↑ CD↓ RCD↓ PSNR↑ SSIM↑ CD↓ RCD↓
25.12 0.5662 28.6160 0.0911 27.16 0.5967 - - 25.50 0.5528 38.5127 0.2149 27.07 0.6417 46.1504 0.2144
25.28 0.6910 6.3486 0.0048 27.52 0.7907 - - 25.75 0.6588 16.8615 0.0102 26.58 0.7488 8.9845 0.0058
23.96 0.6525 5.9112 0.0040 23.43 0.7094 - - 23.46 0.6009 15.0229 0.0099 23.68 0.6510 13.2131 0.0185
26.65 0.7001 5.4482 0.0033 27.98 0.7946 - - 26.21 0.6635 10.4558 0.0051 27.55 0.7724 8.0739 0.0040
@@ 4.1 Problem Formulation
In dynamic driving scenarios, we are given a sequence of radar measurements $S=\{S_{0},S_{1},...,S_{n-1}\}$, where $S_{i}\in\mathbb{R}^{N_{\theta}\times N_{\delta}}$, along with the corresponding radar poses $H_{s}=\{H_{0},H_{1},...,H_{n-1}\}$, where $H_{i}\in\mathrm{SE}(3)$, and timestamps $T_{s}=\{t_{0},t_{1},...,t_{n-1}\}$, where $t_{i}\in\mathbb{R}$. Here, $N_{\theta}$ denotes the number of azimuth directions (beams), and $N_{\delta}$ is the number of range bins per beam. Each bin records a 1D reflected power value measured at the corresponding azimuth and range.
@@ 4.2 RF4D
The latent feature $\chi$ is then passed to $f_{\alpha}$, an MLP that predicts occupancy using a Gumbel-Sigmoid activation [26] at the output. Note that Gumbel-Sigmoid is to encourage the predicted $\alpha$ to approximate binary values, functioning as a soft binary mask aligned with physical interpretation of occupancy: 0 implies empty and 1 implies occupied. Since RCS also depends on the incident angle of radar waveform (view direction), to predict RCS $\sigma$, we concatenate $\chi$ with the encoded direction $\mathbf{d}$ and pass them to $f_{\sigma}$. Here we use spherical harmonics $\mathcal{S}$ [67] to encode $\mathbf{d}$. Finally, we obtain the two predicted radar-specific characteristics:
@@
To ensure temporal consistency in occupancy predictions across frames, we introduce a scene flow module that models the motion of scene points over adjacent frames. Given a latent feature representation $\chi$ at time $t$, this module predicts motion offsets $\Delta x$ through an MLP $f_{\Delta x}$:
$\Delta x=f_{\Delta x}(\chi).$ (4)
The offset $\Delta x\in\mathbb{R}^{6}$ consists of two parts: the first three dimensions represent the offset to the previous frame $\Delta x^{-}$, and the last three dimensions represent the offset to the next frame $\Delta x^{+}$. Each spatial point is then warped to its estimated position in the adjacent frames. The warped points are passed through the feature MLP $f_{\chi}$ and the occupancy MLP $f_{\alpha}$ to estimate the occupancies at adjacent frames:
$\alpha^{t-\Delta t}=f_{\alpha}(f_{\chi}(\mathcal{H}(x+\Delta x^{-}),\mathcal{T}(t-\Delta t))),$ (5)
$\alpha^{t+\Delta t}=f_{\alpha}(f_{\chi}(\mathcal{H}(x+\Delta x^{+}),\mathcal{T}(t+\Delta t))).$
We regularize these occupancy predictions across time to promote temporal coherence. This encourages the model to produce stable and consistent occupancy, especially for moving objects.
@@
Synthesizing radar signals requires incorporating the radar sensing physics and cannot directly apply traditional optical volume rendering in NeRF. The received radar power follows the model in Equation 1, where power depends on the radar cross-section (RCS) $\sigma$ of the target and its distance $\delta$ from the sensor. Following [4], constant factors such as transmitted power and antenna gain are omitted since they remain fixed during training. A base-10 logarithmic transformation is applied to match the decibel scale commonly used in radar data. Instead of disentangling occupancy and reflectance like  [4], we treat the predicted occupancy $\alpha$ as a soft gating mask that activates radar responses only when a target is predicted physically present. This mechanism suppresses noise and reduces the effect of multipath interference. The resulting rendering is given by: